### v1

# UniFlow: Towards Zero-Shot LiDAR Scene Flow for Autonomous Vehicles via Cross-Domain Generalization

Abstract
LiDAR scene flow is the task of estimating per-point 3D motion between consecutive point clouds. Recent methods achieve centimeter-level accuracy on popular autonomous vehicle (AV) datasets, but are typically only trained and evaluated on a single sensor. In this paper, we aim to learn general motion priors that transfer to diverse and unseen LiDAR sensors. However, prior work in LiDAR semantic segmentation and 3D object detection demonstrate that naively training on multiple datasets yields worse performance than single dataset models. Interestingly, we find that this conventional wisdom does not hold for motion estimation, and that state-of-the-art scene flow methods greatly benefit from cross-dataset training. We posit that low-level tasks such as motion estimation may be less sensitive to sensor configuration; indeed, our analysis shows that models trained on fast-moving objects (e.g., from highway datasets) perform well on fast-moving objects, even across different datasets. Informed by our analysis, we propose UniFlow, a family of feedforward models that unifies and trains on multiple large-scale LiDAR scene flow datasets with diverse sensor placements and point cloud densities. Our frustratingly simple solution establishes a new state-of-the-art on Waymo and nuScenes, improving over prior work by 5.1% and 35.2% respectively. Moreover, UniFlow achieves state-of-the-art accuracy on unseen datasets like TruckScenes, outperforming prior TruckScenes-specific models by 30.1%. We include additional visuals on our project page.
RGB |
Sensor |
BEV |
1 Introduction
Recent LiDAR scene flow methods achieve remarkable performance on popular autonomous vehicle (AV) datasets [vedder2024neural, zeroflow, zhang2025deltaflow, khatri2024can] like Argoverse 2, nuScenes, and Waymo [sun2020scalability, nuscenes, Argoverse2_2021]. Notably, current state-of-the-art approaches train and evaluate on each dataset independently and do not attempt to generalize to unseen sensor configurations. Intuitively, since each dataset has different sensor placements and point cloud densities, training separate models for each domain seems reasonable (Figure 1). For example, Argoverse 2 (AV2) [Argoverse2_2021] uses two out-of-phase 32 beam LiDARs, nuScenes [nuscenes] uses one 32 beam sensor, and Waymo [sun2020scalability] uses a custom 64 beam sensor. Indeed, it is notoriously challenging to train multi-dataset models for LiDAR-based 3D object detection [wang2020train, soum2023mdt3d] and semantic segmentation [kim2024rethinking, saltori2023walking] that improve over single dataset models. Interestingly, our analysis challenges this conventional wisdom for LiDAR-based scene flow. Unlike “high-level” tasks like LiDAR semantic segmentation or 3D object detection that require human-defined taxonomies [lambert2020mseg], we argue that LiDAR scene flow models capture universal “low-level” geometric features that should, in principle, generalize better across sensors.
Cross-Domain Generalization for Low-Level Vision. Notably, we observe similar trends in RGB-based low-level vision tasks, where models learn generalizable geometric representations that transfer well across datasets. For example, FlowNet [dosovitskiy2015flownet, ilg2017flownet] and RAFT [teed2020raft] show that optical flow models trained on synthetic datasets (e.g. FlyingChairs, FlyingThings3D, Sintel) generalize surprisingly well to casually captured videos. More recently, zero-shot monocular depth estimators like MoGe [wang2025moge] and monocular scene flow estimators like ZeroMSF [liang2025zero] demonstrate that training on diverse datasets significantly improves zero-shot performance. Therefore, we posit that low-level LiDAR-based scene flow may similarly generalize to different sensors if trained at scale.
To better understand the problem, we first train dataset-specific models for AV2 [Argoverse2_2021], Waymo [sun2020scalability] and nuScenes [nuscenes] and evaluate zero-shot cross-domain performance. Surprisingly, Flow4D [kim2024flow4d] trained on Waymo achieves similar Dynamic Mean EPE on AV2 as Flow4D trained on AV2. Importantly, Waymo’s 64 beam LiDAR has considerably more point density than AV2’s two out-of-phase 32-beam LiDARs (Fig. 1), indicating that point density does not significantly impact LiDAR scene flow generalization. Further, Flow4D (Waymo) achieves lower EPE (i.e. better performance) than Flow4D (AV2) on fast moving objects in AV2 since there are more fast movers in Waymo than in AV2 (Table 1). This suggests that the accuracy of motion prediction is highly correlated with object velocity (Figure 2); in order to accurately predict fast motion, one should train on fast motion. This suggests that the key to generalization is training on a diverse combination of datasets with a wide range of object velocities.
| Train | Test | FD (cm) | Dyn. Mean | [0, 0.5) | [0.5, 1.0) | [1.0, 2.0) | [2.0, ) |
| AV2 | AV2 | 8.55 | 0.192 | 0.020 | 0.119 | 0.114 | 0.103 |
| Waymo | 8.31 | 0.236 | 0.016 | 0.077 | 0.079 | 0.350 | |
| nuScenes | 20.40 | 0.463 | 0.017 | 0.249 | 0.334 | 0.610 | |
| Waymo | AV2 | 8.90 | 0.199 | 0.021 | 0.119 | 0.125 | 0.098 |
| Waymo | 4.58 | 0.215 | 0.007 | 0.064 | 0.051 | 0.130 | |
| nuScenes | 16.04 | 0.408 | 0.017 | 0.217 | 0.319 | 0.540 | |
| nuScenes | AV2 | 16.08 | 0.357 | 0.039 | 0.239 | 0.272 | 0.395 |
| Waymo | 17.24 | 0.364 | 0.022 | 0.222 | 0.209 | 0.598 | |
| nuScenes | 7.97 | 0.230 | 0.018 | 0.100 | 0.109 | 0.370 |
Towards Zero-Shot LiDAR Scene Flow. We argue that scaling up LiDAR-based scene flow methods will be a key enabler for 3D motion understanding and dynamic reconstruction in diverse environments. To this end, we retrain state-of-the-art methods on supervised data from AV2, Waymo, and nuScenes and demonstrate significant improvements for both in-distribution and out-of-distribution generalization. We denote models trained with multiple datasets as UniFlow. To the best of our knowledge, UniFlow is the first to demonstrate the benefits of cross-domain training for LiDAR scene flow. Further, UniFlow models demonstrate remarkable zero-shot accuracy on TruckScenes [fent2024man], outperforming prior dataset-specific models by 30.1% (Table 6). Lastly, we find that cross-domain training on near-range (i.e 35 meters from the ego-vehicle) points improves scene flow generalization to unseen far-range points (Table 7). Intuitively, sparse LiDAR points near the ego-vehicle mimic the distribution of dense LiDAR points far away from the vehicle [peri2023empirical].
Why Does Multi-Dataset Training Work? Our experiments suggest that multi-dataset training improves LiDAR scene flow performance because scene flow is inherently class-agnostic. This allows us to avoid complications arising from dataset-specific label definitions (Table 8). For example, nuScenes defines bicycle as excluding the rider, while Waymo includes the rider [madan2023revisiting, robicheaux2025roboflow100]. Such label ambiguity poses a significant challenge for semantic tasks, potentially explaining why prior LiDAR-based semantic tasks like 3D object detection and semantic segmentation do not benefit from training on multiple datasets. Furthermore, we find that with careful augmentations, we can simulate diverse sensor configurations that improve cross-dataset generalization (Figure 4). For example, dropping out LiDAR beams from a Waymo point cloud makes the data look more like a point cloud from AV2 or nuScenes, while augmenting the height of a point cloud makes a sensor mounted on a sedan look like it was mounted on a truck.
Contributions. We present three major contributions. First, we highlight that dataset-specific scene flow models already show signs of cross-domain generalization, motivating our investigation into multi-dataset training. Next, we demonstrate that re-training existing scene flow methods on diverse, large-scale data yields state-of-the-art performance on nuScenes and Waymo, improving over our baselines by 35.2% and 5.1% respectively. Finally, we show that UniFlow outperforms dataset-specific models by 30.1% on TruckScenes, challenging conventional wisdom on LiDAR-based cross-domain generalization.
2 Related Work
Scene Flow Estimation is the task of describing the 3D motion field between temporally successive point clouds [vedula2005three, khatri2024can, zhang2025deltaflow], and is an important primitive for 3D motion understanding and dynamic scene reconstruction [wang2025fluxd]. Early approaches [wei2020pv, scoop, wang2023dpvraft, zhang2024gmsf] learned point-wise features to estimate per-point flow but struggled to scale to large outdoor environments [zhang2024deflow, zeroflow]. More recent work [khoche2025ssf, kim2024flow4d, mambaflow, hoffmann2025floxels, zeroflow] jointly estimates flow for all points in a scene. Contemporary methods can be broadly classified into feedforward models and optimization-based methods. Feedforward models directly learn a mapping between point cloud pairs and flow fields, but require large-scale human annotations [fastflow3d, kim2024flow4d]. For real world datasets (typically from the autonomous vehicle domain), these human annotations are provided in the form of 3D bounding boxes and tracks for every object in the scene. In contrast, optimization-based methods do not require labeled data, and instead optimize a learned representation per-scene [li2021neural, li2023fast]. This per-scene optimization is prohibitively expensive, making it difficult to scale to large datasets. For example, EulerFlow [vedder2024neural] achieves state-of-the-art unsupervised accuracy on many popular AV datasets, but takes over 24 hours to optimize each scene. In summary, feedforward models achieve efficient inference but demonstrate limited generalization beyond their training data, while optimization-based methods produce high quality flow for diverse scenes, but are too slow for real-time applications. Our work aims to reconcile this trade-off by training a single feedforward model that achieves robust generalization across diverse datasets and sensors.
Cross-Domain Generalization for LiDAR is a long-standing challenge in LiDAR-based perception. Prior works in 3D object detection [malic2025gblobs, hegde2025multimodal] and semantic segmentation [kim2024rethinking, caunes2025seg_3d_by_pc2d, xu2025superflowpp] show that models trained on multiple datasets often perform worse than dataset-specific models. This performance drop can be largely attributed to diverse sensor hardware (e.g., number of beams, scan pattern, point cloud density), environmental conditions (e.g., weather, geography), and sensor placement across datasets. To address these challenges, several methods have introduced data augmentation strategies [sun2024empirical], such as random point dropping [wang2021pointaugmenting] and object scaling [cen2022open], while others [Liu_2024_CVPR, michele2024saluda] have proposed unsupervised domain adaptation techniques that align feature distributions across datasets. However, such data augmentation strategies primarily address the geometric domain gap between datasets. For motion-centric tasks such as scene flow, we argue that understanding the distribution of object velocities between datasets represents a critical yet overlooked axis for addressing cross-domain generalization. To this end, our work presents the first systematic study of this velocity domain gap.
3 UniFlow: Towards Zero-Shot LiDAR Scene Flow for Autonomous Vehicles
UniFlow aims to learn general motion priors that transfer to diverse and unseen LiDAR sensors by training a single scene flow model across diverse autonomous driving datasets. In this section, we describe the process of unifying four popular AV datasets and our training strategy.
Dataset Unification. We unify four widely used AV datasets to facilitate multi-dataset training. Importantly, we standardize LiDAR sensor frame rates and annotation frequencies to ensure that the time interval between LiDAR sweeps is consistent across datasets. This step is crucial because state-of-the-art scene flow methods parameterize motion as the displacement between two consecutive LiDAR frames. Specifically, we down-sample high-frequency datasets like nuScenes (which provides LiDAR sweeps at 20 Hz) to 10 Hz to match AV2 and Waymo. For datasets with sparse scene flow annotations (e.g., nuScenes annotates tracks at 2 Hz), we train solely on the annotated frames and their real adjacent scans. This ensures that we do not introduce any label noise or motion artifacts due to interpolation. Lastly, we apply LineFit [himmelsbach2010fast] for ground point removal for all datasets.
Unified Multi-Dataset Training. We retrain existing state-of-the-art scene flow architectures (SSF [khoche2025ssf], Flow4D [kim2024flow4d], and Flow [zhang2025deltaflow]) on the unified dataset mixture of nuScenes, AV2, and Waymo. Importantly, we do not re-weight dataset frequencies or apply dataset-specific augmentations. Instead, we employ a minimal yet effective augmentation set including height randomization to account for sensor placement and LiDAR ray dropping to simulate varying LiDAR point densities. This unified training paradigm encourages models to learn domain-agnostic geometric priors rather than memorizing dataset-specific motion patterns, resulting in significant zero-shot generalization improvements.
| Dataset | Split | # Scenes | # Frames | LiDAR Setup | Ego-Vehicle | Scenario |
| Argoverse 2 | Train | 700 | 110,071 | 2 x 32-beam | Car | Urban / City |
| Val | 150 | 23,547 | ||||
| Waymo | Train | 798 | 155,000 | 1 x 64-beam | Car | Urban / Suburban |
| Val | 202 | 40,000 | ||||
| nuScenes | Train | 700 | 137,575 / 27,392 | 1 x 32-beam | Car | Urban / City |
| Val | 150 | 29,126 / 5,798 | ||||
| TruckScenes | Train | 524 | 101,902 / 20,380 | 2 x 64-beam | Truck | Highway |
| Val | 75 | 14,625 / 2,925 |
Overall Performance |
Rare Vehicles |
Long Range ( m) |
(a) Flow
(b) Flow (UniFlow)
(c) Ground truth
4 Experiments
Datasets. We train and evaluate UniFlow on four large-scale autonomous driving datasets to rigorously evaluate cross-domain generalization to diverse sensor configurations, ego-vehicle platforms, and driving scenarios (Table 2). AV2 and nuScenes collect data in dense urban environments with 32-beam LiDARs. Waymo uses a custom LiDAR sensor to collect denser point clouds from urban and suburban environments with more varied driving dynamics. We evaluate zero-shot generalization using TruckScenes [fent2024man] since it collects data in different environments (high-speed highway scenes vs. urban and suburban scenes), with different sensors (32-beam LiDARs vs. 64-beam LiDARs), using a different vehicle type (truck vs. sedan). These four datasets allow us to systematically analyze the impact of point density (32-beam vs. 64-beam), ego-vehicle perspective (car vs. truck), and velocity (urban vs. highway) on cross-domain generalization.
Metrics. We use Dynamic Bucket-Normalized EPE [khatri2024can] as our primary metric to evaluate cross-domain generalization, particularly across diverse velocity distributions. This protocol is specifically designed to be both class-aware and speed-normalized. By normalizing motion estimation error by an object’s speed, Dynamic Bucket-Normalized EPE measures the fraction of motion not described, enabling fair comparison between slow-moving pedestrians and fast-moving vehicles. We also report Three-Way EPE [Chodosh_2024_WACV] for completeness, but note that it is a biased estimator of performance.
Baselines. We compare UniFlow against both self-supervised and fully-supervised methods: NSFP [li2021neural], FastNSF [li2023fast], ICPFlow [lin2024icp], SeFlow [zhang2024seflow], SeFlow++ [zhang2025himo], EulerFlow [vedder2024neural], VoteFlow [lin2025voteflow], TrackFlow [khatri2024can], DeFlow [zhang2024deflow], SSF [khoche2025ssf], Flow4D [kim2024flow4d] and Flow [zhang2025deltaflow]. To ensure fairness, Argoverse 2 test-set results are obtained directly from the public leaderboard. We use Waymo, nuScenes and TruckScenes baselines as implemented in OpenSceneFlow111https://github.com/KTH-RPL/OpenSceneFlow with standard training hyperparameters.
| Methods | Three-way EPE (cm) | Dynamic Bucket-Normalized | |||||||
| Mean | FD | FS | BS | Mean | Car | Other | Pedestr. | VRU | |
| Unsupervised | |||||||||
| NSFP | 6.06 | 11.58 | 3.16 | 3.44 | 0.422 | 0.251 | 0.331 | 0.722 | 0.383 |
| FastNSF | 11.18 | 16.34 | 8.14 | 9.07 | 0.383 | 0.296 | 0.413 | 0.500 | 0.322 |
| SeFlow | 4.86 | 12.14 | 1.84 | 0.60 | 0.309 | 0.214 | 0.291 | 0.464 | 0.265 |
| ICP Flow | 6.50 | 13.69 | 3.32 | 2.50 | 0.331 | 0.195 | 0.331 | 0.435 | 0.363 |
| EulerFlow | 4.23 | 4.98 | 2.45 | 5.25 | 0.130 | 0.093 | 0.141 | 0.195 | 0.093 |
| Supervised | |||||||||
| TrackFlow | 4.73 | 10.30 | 3.65 | 0.24 | 0.269 | 0.182 | 0.305 | 0.358 | 0.230 |
| DeFlow | 3.43 | 7.32 | 2.51 | 0.46 | 0.276 | 0.113 | 0.228 | 0.496 | 0.266 |
| SSF | 2.73 | 5.72 | 1.76 | 0.72 | 0.181 | 0.099 | 0.162 | 0.292 | 0.169 |
| Flow4D | 2.24 | 4.94 | 1.31 | 0.47 | 0.145 | 0.087 | 0.150 | 0.216 | 0.127 |
| Flow | 2.11 | 4.33 | 1.37 | 0.64 | 0.113 | 0.077 | 0.129 | 0.149 | 0.096 |
| SSF (UniFlow, Ours) | 2.23 | 4.89 | 1.31 | 0.51 | 0.156 | 0.102 | 0.147 | 0.241 | 0.134 |
| Flow4D (UniFlow, Ours) | 2.07 | 4.50 | 1.31 | 0.40 | 0.132 | 0.072 | 0.131 | 0.218 | 0.107 |
| Flow (UniFlow, Ours) | 2.07 | 4.41 | 1.26 | 0.54 | 0.118 | 0.073 | 0.136 | 0.164 | 0.098 |
| Methods | Three-way EPE (cm) | Dynamic Bucket-Normalized | ||||||
| Mean | FD | FS | BS | Mean | Car | Pedestr. | VRU | |
| Unsupervised | ||||||||
| NSFP | 10.05 | 17.12 | 10.81 | 2.21 | 0.574 | 0.315 | 0.823 | 0.584 |
| ZeroFlow | 8.64 | 22.40 | 1.57 | 1.96 | 0.770 | 0.444 | 0.982 | 0.884 |
| SeFlow | 8.50 | 20.81 | 2.14 | 2.57 | 0.328 | 0.305 | 0.485 | 0.185 |
| SeFlow++ | 3.63 | 9.30 | 0.87 | 0.71 | 0.232 | 0.201 | 0.521 | 0.247 |
| Supervised | ||||||||
| DeFlow | 4.46 | 9.80 | 2.59 | 0.98 | - | - | - | - |
| SSF | 2.19 | 5.62 | 0.74 | 0.23 | 0.264 | 0.097 | 0.469 | 0.225 |
| Flow4D | 1.82 | 4.58 | 0.61 | 0.28 | 0.215 | 0.074 | 0.423 | 0.146 |
| Flow | 1.40 | 3.27 | 0.58 | 0.35 | 0.198 | 0.073 | 0.407 | 0.114 |
| SSF (UniFlow, Ours) | 1.85 | 4.74 | 0.58 | 0.22 | 0.234 | 0.075 | 0.434 | 0.195 |
| Flow4D (UniFlow, Ours) | 1.60 | 3.83 | 0.68 | 0.30 | 0.191 | 0.064 | 0.400 | 0.110 |
| Flow (UniFlow, Ours) | 1.38 | 3.07 | 0.61 | 0.46 | 0.188 | 0.065 | 0.403 | 0.096 |
| Methods | Three-way EPE (cm) | Dynamic Bucket-Normalized | |||||||
| Mean | FD | FS | BS | Mean | Car | Other | Pedestr. | VRU | |
| Unsupervised | |||||||||
| NSFP | 10.79 | 20.26 | 4.88 | 7.23 | 0.602 | 0.463 | 0.456 | 0.829 | 0.662 |
| FastNSF | 12.16 | 18.20 | 6.11 | 12.18 | 0.560 | 0.436 | 0.523 | 0.737 | 0.543 |
| SeFlow | 8.19 | 16.15 | 3.97 | 4.45 | 0.554 | 0.396 | 0.635 | 0.726 | 0.419 |
| Supervised | |||||||||
| DeFlow | 3.98 | 6.99 | 3.45 | 1.50 | 0.314 | 0.163 | 0.286 | 0.533 | 0.275 |
| SSF | 3.00 | 6.55 | 2.04 | 0.41 | 0.220 | 0.142 | 0.197 | 0.398 | 0.144 |
| Flow4D | 3.46 | 7.97 | 1.85 | 0.55 | 0.230 | 0.160 | 0.241 | 0.345 | 0.176 |
| Flow | 2.33 | 4.83 | 1.37 | 0.79 | 0.216 | 0.138 | 0.219 | 0.327 | 0.181 |
| SSF (UniFlow, Ours) | 1.97 | 4.33 | 1.38 | 0.20 | 0.144 | 0.081 | 0.131 | 0.267 | 0.097 |
| Flow4D (UniFlow, Ours) | 3.01 | 7.28 | 1.47 | 0.28 | 0.196 | 0.137 | 0.219 | 0.272 | 0.157 |
| Flow (UniFlow, Ours) | 1.63 | 3.19 | 1.29 | 0.42 | 0.140 | 0.098 | 0.137 | 0.232 | 0.091 |
| Methods | Three-way EPE (cm) | Dynamic Bucket-Normalized | |||||||
| Mean | FD | FS | BS | Mean | Car | Other | Pedestr. | VRU | |
| Unsupervised | |||||||||
| NSFP | 45.63 | 120.45 | 2.07 | 14.38 | 0.658 | 0.303 | 0.350 | 1.221 | 0.758 |
| FastNSF | 30.72 | 59.44 | 3.35 | 29.38 | 0.588 | 0.218 | 0.376 | 1.124 | 0.635 |
| SeFlow | 37.41 | 103.97 | 1.56 | 6.68 | 0.681 | 0.494 | 0.752 | 0.886 | 0.591 |
| ICP Flow | 58.82 | 169.91 | 1.43 | 5.12 | 0.472 | 0.302 | 0.614 | 0.596 | 0.376 |
| Supervised | |||||||||
| DeFlow | 7.30 | 16.47 | 1.67 | 3.77 | 0.570 | 0.180 | 0.410 | 0.970 | 0.730 |
| SSF | 6.20 | 15.17 | 1.01 | 2.44 | 0.453 | 0.215 | 0.308 | 0.875 | 0.414 |
| Flow4D | 16.14 | 44.87 | 1.71 | 1.85 | 0.456 | 0.176 | 0.351 | 0.885 | 0.413 |
| Flow | 7.28 | 16.26 | 1.36 | 4.52 | 0.402 | 0.196 | 0.400 | 0.690 | 0.323 |
| Zero-shot | |||||||||
| SSF (UniFlow, Ours) | 35.23 | 103.72 | 1.69 | 0.27 | 0.435 | 0.149 | 0.455 | 0.669 | 0.466 |
| Flow4D (UniFlow, Ours) | 23.59 | 68.41 | 1.78 | 0.57 | 0.281 | 0.088 | 0.277 | 0.530 | 0.230 |
| Flow (UniFlow, Ours) | 15.76 | 45.76 | 1.02 | 1.32 | 0.283 | 0.101 | 0.293 | 0.513 | 0.226 |
Comparison to State-of-the-Art Methods. We demonstrate the effectiveness of the UniFlow framework by retraining SSF, Flow4D, and Flow on multiple datasets and compare their performance against recent approaches on popular scene flow benchmarks (Tables 3, 4, and 5). On the Argoverse 2 test set, Flow4D (UniFlow) improves over Flow4D by 8.97% and SSF (UniFlow) improves over SSF by 13.81% in Dynamic Mean EPE. Moreover, Flow4D (UniFlow) achieves near-parity with EulerFlow at a fraction of the runtime. We see similar improvements between Flow4D and SSF and their UniFlow counterparts on the Waymo (Table 4) and nuScenes (Table 5) validation sets. Interestingly, SSF (UniFlow) outperforms Flow4D (UniFlow) on nuScenes, suggesting that the SSF architecture is naturally better suited for processing sparse point clouds.
TruckScenes Zero-Shot Performance. We demonstrate UniFlow’s strong zero-shot generalization capabilities with extensive evaluations on TruckScenes (Table 6). All baseline methods (except UniFlow) were trained on TruckScenes. Despite never being explicitly trained on high speed highway scenes, Flow4D (UniFlow) significantly outperforms the dataset-specific Flow4D on Dynamic Mean EPE. Notably, Flow4D (UniFlow) was only trained on urban driving scenes collected from a sedan, indicating that our approach can successfully generalize across significant domain gaps. Table 6 also highlight the importance of the speed normalized and bucketed EPE metric [khatri2024can], as Threeway-EPE results don’t adequately measure Flow4D’s (UniFlow) significant improvements on pedestrian and VRU performance.
| Dataset | Method | Mean | 0–35 | 35–50 | 50–75 | 75–100 | 100–inf |
| AV2 | Flow | 0.526 | 0.117 | 0.492 | 0.745 | 0.708 | 0.568 |
| FlowLR | 0.259 | 0.117 | 0.205 | 0.240 | 0.325 | 0.409 | |
| FlowLR (UniFlow, Ours) | 0.226↓12.7% | 0.114↓2.6% | 0.198↓3.4% | 0.219↓8.8% | 0.260↓20.0% | 0.339↓17.1% | |
| Waymo | Flow | 0.379 | 0.055 | 0.455 | 0.628 | - | - |
| FlowLR | 0.115 | 0.068 | 0.112 | 0.166 | - | - | |
| FlowLR (UniFlow, Ours) | 0.105↓9.1% | 0.063↓7.4% | 0.104↓7.1% | 0.149↓10.2% | - | - | |
| nuScenes | Flow | 0.389 | 0.172 | 0.380 | 0.633 | 0.606 | - |
| FlowLR | 0.267 | 0.181 | 0.242 | 0.318 | 0.447 | - | |
| FlowLR (UniFlow, Ours) | 0.248↓7.1% | 0.166↓7.9% | 0.209↓13.6% | 0.301↓5.3% | 0.405↓9.4% | - | |
| TruckScenes | Flow | 1.679 | 1.211 | 1.514 | 1.901 | 1.835 | 1.936 |
| FlowLR | 1.479 | 1.184 | 1.311 | 1.487 | 1.582 | 1.831 | |
| FlowLR,ZS (UniFlow, Ours) | 0.880↓40.6% | 0.786↓33.6% | 0.826↓37.0% | 0.694↓53.3% | 0.779↓50.8% | 1.317↓28.1% |
Analysis of Long-Range Generalization. To evaluate whether unified training produces range-robust representations, we evaluate how models trained on short-range LiDAR data transfer to longer-range settings without retraining. Specifically, we extend the voxelization range during inference while keeping the voxel size identical to the original trained model. This change alters only the spatial coverage of the dynamic voxelization module, requiring a simple weight transfer step to adapt to the new grid dimensions while preserving all learned features. As shown in Table 7, UniFlowLR consistently outperforms FlowLR across all distance intervals on AV2, reducing Dynamic Mean EPE by 12.7% on average and up to 20.0% in the 75–100 m range. This indicates that unified multi-dataset training produces representations that remain robust even when the perception range is significantly expanded. On TruckScenes, UniFlowLR,ZS achieves a remarkable 40.6% reduction in overall error compared to FlowLR, with improvements exceeding 50% in the 50–100 m range, despite never being trained on this dataset. These results highlight that the representations learned through unified multi-dataset training naturally scale to larger spatial extents.
4.1 Ablation Studies
Analysis of Semantic Cross-Domain Generalization. We evaluate the benefits unified training on both geometric and semantic understanding by extending Flow4D with a multi-task semantic head. This new Flow4D variant simultaneously predicts scene flow and per-point semantics using the same backbone features. Since all datasets follow different taxonomies, we map their annotations to four coarse categories (e.g. car, other, pedestrian, and VRU) and jointly train the model with an additional class-balanced cross-entropy loss. As shown in Table 8, adding the semantic head has almost no effect on scene flow accuracy, indicating that the auxiliary task does not interfere with motion estimation. Consistent with our original findings, unified multi-dataset training continues to improve both in-domain and zero-shot performance for scene flow. However, training the semantic prediction head on the unified dataset yields mixed results. Models trained on unified datasets show small in-domain gains, but perform substantially worse than dataset-specific model when evaluated on unseen domains. This supports our hypothesis that dataset unification primarily benefits low-level geometric representations, while high-level semantics remain sensitive to dataset-specific biases.
| Dataset | Scene Flow | Semantic Classification | |||||||
| Train | Test | 3-Way (cm) | Dyn. Norm. | mIoU | Accuracy | CAR | OTHER | PED | VRU |
| AV2 | AV2 | 3.46 | 0.204 | 75.3% | 93.9% | 90.9% | 86.2% | 88.2% | 36.0% |
| Waymo | Waymo | 1.74 | 0.202 | 79.1% | 98.9% | 99.0% | - | 87.1% | 51.3% |
| nuScenes | nuScenes | 2.99 | 0.214 | 70.5% | 89.8% | 82.8% | 83.3% | 80.5% | 35.4% |
| TruckScenes | TruckScenes | 15.56 | 0.462 | 64.6% | 95.3% | 91.0% | 92.9% | 32.6% | 42.1s% |
| Unified | AV2 | 3.26 | 0.178 | 78.8% | 94.2% | 91.4% | 85.8% | 90.0% | 47.9% |
| Unified | Waymo | 1.57 | 0.190 | 64.0% | 99.4% | 99.4% | 0.0% | 93.1% | 63.5% |
| Unified | nuScenes | 3.19 | 0.215 | 79.9% | 93.4% | 88.0% | 89.1% | 82.3% | 60.3% |
| Unified | TruckScenes | 24.69 | 0.288 | 40.5% | 66.1% | 53.5% | 46.4% | 20.1% | 41.9% |
Analysis of Training Recipe. We ablate the impact of multi-dataset training, data augmentation, and model scaling on zero-shot TruckScenes performance in Table 9. All baselines are trained on Waymo only to ensure zero-shot evaluation. All model architectures show clear gains from multi-dataset training. Notably, despite training on a dataset three times larger, data augmentation continues to show substantial improvements, indicating that further scaling may yield additional performance gains.
| Method | FD (cm) | Dynamic EPE | [0, 0.5) | [0.5, 1.0) | [1.0, 2.0) | [2.0, ) |
| Flow4D | 116.01 | 0.336 | 0.071 | 0.129 | 0.215 | 0.586 |
| + Unified Dataset | 93.76 ↓22.25 | 0.310 ↓0.026 | 0.040 ↓0.031 | 0.116 ↓0.013 | 0.170 ↓0.045 | 0.594 ↑0.008 |
| + Augmentation | 68.32 ↓25.44 | 0.301 ↓0.009 | 0.047 ↑0.007 | 0.111 ↓0.005 | 0.141 ↓0.029 | 0.407 ↓0.187 |
| + XL Backbone | 68.41 ↑0.09 | 0.281 ↓0.020 | 0.033 ↓0.014 | 0.097 ↓0.014 | 0.119 ↓0.022 | 0.389 ↓0.018 |
| SSF | 170.65 | 0.737 | 0.049 | 0.555 | 0.740 | 0.971 |
| + Unified Dataset | 133.04 ↓37.61 | 0.577 ↓0.160 | 0.042 ↓0.007 | 0.257 ↓0.298 | 0.514 ↓0.226 | 0.827 ↓0.144 |
| + Augmentation | 103.72 ↓29.32 | 0.435 ↓0.142 | 0.036 ↓0.006 | 0.149 ↓0.108 | 0.285 ↓0.229 | 0.637 ↓0.190 |
| Flow | 117.83 | 0.357 | 0.067 | 0.114 | 0.302 | 0.714 |
| + Unified Dataset | 84.16 ↓33.67 | 0.331 ↓0.026 | 0.036 ↓0.031 | 0.119 ↑0.005 | 0.268 ↓0.034 | 0.659 ↓0.055 |
| + Augmentation | 15.76 ↓68.40 | 0.283 ↓0.048 | 0.039 ↑0.003 | 0.094 ↓0.025 | 0.127 ↓0.141 | 0.429 ↓0.230 |
Analysis on Dataset Weighting Strategies. We further examine the impact of per-dataset sampling frequencies on our unified training approach. All models are trained on the unified dataset for a fixed number of iterations, but we vary the per-domain sampling frequency to control how often samples from each distribution are drawn. As shown in Table 10, unified training consistently outperforms the single-dataset baselines, even under extreme weighting schemes. For example, on nuScenes, models trained with only 5% exposure to the target domain and 90% exposure to a very different distribution (Waymo) still surpass the in-domain baseline by 13.5% on Dynamic Mean EPE, indicating that even limited cross-domain exposure provides considerable benefits. Unsurprisingly, one can maximize in-domain performance by slightly biasing the model twoards in-domain data. Lastly, we find that uniformly sampling all datasets with equal probability yields the best out-of-distribution performance on TruckScenes.
| Weight Strategy | Dataset Weights | Mean | FD | FS | BS | Dynamic M |
| TruckScenes | ||||||
| Proportional | 60%W, 25%A, 15%N | 23.59 | 68.41 | 1.78 | 0.57 | 0.281 |
| Uniform | 33%W, 33%A, 33%N | 30.44 | 89.01 | 1.73 | 0.58 | 0.307 |
| Waymo Heavy | 90%W, 5%A, 5%N | 33.25 | 97.48 | 1.72 | 0.56 | 0.312 |
| AV2 Heavy | 90%A, 5%W, 5%N | 29.82 | 87.08 | 1.70 | 0.67 | 0.310 |
| nuScenes Heavy | 90%N, 5%A, 5%W | 31.26 | 91.11 | 1.70 | 0.96 | 0.332 |
| Argoverse 2 | ||||||
| Proportional | 60%W, 25%A, 15%N | 3.24 | 7.83 | 1.26 | 0.63 | 0.176 |
| Uniform | 33%W, 33%A, 33%N | 3.22 | 7.79 | 1.25 | 0.63 | 0.176 |
| Waymo Heavy | 90%W, 5%A, 5%N | 3.28 | 7.87 | 1.28 | 0.69 | 0.189 |
| AV2 Heavy | 90%A, 5%W, 5%N | 3.21 | 7.80 | 1.27 | 0.56 | 0.183 |
| nuScenes Heavy | 90%N, 5%A, 5%W | 3.51 | 8.42 | 1.37 | 0.75 | 0.192 |
| AV2-only | 100%A | 3.62 | 8.55 | 1.62 | 0.69 | 0.192 |
| nuScenes | ||||||
| Proportional | 60%W, 25%A, 15%N | 3.01 | 7.28 | 1.47 | 0.28 | 0.196 |
| Uniform | 33%W, 33%A, 33%N | 2.07 | 4.70 | 1.34 | 0.16 | 0.134 |
| Waymo Heavy | 90%W, 5%A, 5%N | 2.87 | 6.88 | 1.48 | 0.25 | 0.199 |
| AV2 Heavy | 90%A, 5%W, 5%N | 2.93 | 6.93 | 1.55 | 0.30 | 0.222 |
| nuScenes Heavy | 90%N, 5%A, 5%W | 1.75 | 3.81 | 1.29 | 0.15 | 0.109 |
| nuScenes-only | 100%N | 3.46 | 7.97 | 1.85 | 0.55 | 0.230 |
| Waymo | ||||||
| Proportional | 60%W, 25%A, 15%N | 1.60 | 3.83 | 0.68 | 0.30 | 0.191 |
| Uniform | 33%W, 33%A, 33%N | 1.59 | 3.89 | 0.61 | 0.27 | 0.195 |
| Waymo Heavy | 90%W, 5%A, 5%N | 1.50 | 3.75 | 0.55 | 0.21 | 0.203 |
| AV2 Heavy | 90%A, 5%W, 5%N | 1.81 | 4.41 | 0.70 | 0.32 | 0.213 |
| nuScenes Heavy | 90%N, 5%A, 5%W | 1.90 | 4.76 | 0.61 | 0.32 | 0.215 |
| Waymo-only | 100%W | 1.82 | 4.58 | 0.61 | 0.28 | 0.215 |
Analysis on Scaling Laws. We evaluate Flow4D (UniFlow) with varying amounts of training data on both in-distribution benchmarks (AV2, nuScenes, and Waymo) and an out-of-distribution benchmark (TruckScenes) in Fig 4. As expected, larger training sets reduce Dynamic Mean EPE. However, data augmentation proves far more critical for out-of-distribution performance, while its effect on in-distribution performance remains minimal.
| Method | Three-way EPE (cm) | Dynamic Bucket-Normalized | |||||||
| Mean | FD | FS | BS | Dyn. Mean | Car | Other | Pedestr. | VRU | |
| 2 Hz | |||||||||
| Flow4D | 30.79 | 88.10 | 3.26 | 1.02 | 0.495 | 0.364 | 0.494 | 0.613 | 0.509 |
| SSF | 31.16 | 89.82 | 3.07 | 0.59 | 0.520 | 0.363 | 0.498 | 0.684 | 0.535 |
| Flow | 26.41 | 74.43 | 3.12 | 1.66 | 0.457 | 0.370 | 0.471 | 0.545 | 0.440 |
| Flow4D (UniFlow, Ours) | 27.52 | 79.54 | 2.36 | 0.65 | 0.444 | 0.297 | 0.415 | 0.587 | 0.476 |
| SSF (UniFlow, Ours) | 28.67 | 83.11 | 2.50 | 0.39 | 0.503 | 0.296 | 0.432 | 0.671 | 0.611 |
| Flow (UniFlow, Ours) | 23.87 | 68.26 | 2.46 | 0.88 | 0.398 | 0.306 | 0.420 | 0.494 | 0.494 |
| 5 Hz | |||||||||
| Flow4D | 9.58 | 25.64 | 2.45 | 0.65 | 0.323 | 0.215 | 0.312 | 0.432 | 0.333 |
| SSF | 8.74 | 23.02 | 2.74 | 0.47 | 0.311 | 0.179 | 0.286 | 0.489 | 0.291 |
| Flow | 6.43 | 16.54 | 1.98 | 0.76 | 0.264 | 0.173 | 0.271 | 0.369 | 0.242 |
| Flow4D (UniFlow, Ours) | 8.02 | 21.78 | 1.92 | 0.37 | 0.290 | 0.190 | 0.265 | 0.405 | 0.302 |
| SSF (UniFlow, Ours) | 6.75 | 17.96 | 2.06 | 0.22 | 0.285 | 0.135 | 0.218 | 0.454 | 0.333 |
| Flow (UniFlow, Ours) | 5.26 | 13.44 | 1.81 | 0.58 | 0.232 | 0.150 | 0.220 | 0.343 | 0.214 |
| 10 Hz (Standard Frame Rate) | |||||||||
| Flow4D | 3.46 | 7.97 | 1.85 | 0.55 | 0.230 | 0.160 | 0.241 | 0.241 | 0.176 |
| SSF | 3.00 | 6.55 | 2.04 | 0.41 | 0.220 | 0.142 | 0.197 | 0.398 | 0.144 |
| Flow | 2.05 | 4.19 | 1.36 | 0.59 | 0.173 | 0.132 | 0.184 | 0.261 | 0.117 |
| Flow4D (UniFlow, Ours) | 3.01 | 7.28 | 1.47 | 0.28 | 0.196 | 0.137 | 0.219 | 0.272 | 0.157 |
| SSF (UniFlow, Ours) | 1.97 | 4.33 | 1.38 | 0.20 | 0.144 | 0.081 | 0.131 | 0.267 | 0.097 |
| Flow (UniFlow, Ours) | 1.63 | 3.18 | 1.29 | 0.42 | 0.140 | 0.098 | 0.137 | 0.232 | 0.091 |
| 20 Hz | |||||||||
| Flow4D | 2.55 | 5.79 | 1.35 | 0.50 | 0.272 | 0.196 | 0.315 | 0.377 | 0.200 |
| SSF | 2.94 | 6.72 | 1.73 | 0.38 | 0.316 | 0.177 | 0.303 | 0.533 | 0.250 |
| Flow | 1.60 | 3.28 | 0.48 | 0.48 | 0.220 | 0.167 | 0.231 | 0.327 | 0.156 |
| Flow4D (UniFlow, Ours) | 1.95 | 4.65 | 1.00 | 0.21 | 0.230 | 0.171 | 0.269 | 0.315 | 0.163 |
| SSF (UniFlow, Ours) | 2.13 | 5.16 | 1.11 | 0.12 | 0.246 | 0.127 | 0.275 | 0.395 | 0.189 |
| Flow (UniFlow, Ours) | 1.30 | 2.60 | 0.96 | 0.34 | 0.010 | 0.153 | 0.215 | 0.277 | 0.138 |
Analysis on Frame-Rate. We evaluate the generalization of Flow4D (UniFlow), SSF (UniFlow) and Flow (UniFlow) across different frame rates in Table 11. Slower frame rates approximate a slower-moving ego-vehicle, while faster frame rates approximate a faster one. Although all models are trained using 10 Hz annotations, multi-dataset training consistently improves performance, especially at slower frame rates. Interestingly, higher frame rates have a smaller negative effect on performance, likely because models only need to estimate smaller displacement vectors for each timestep.
Limitations and Future Work. We investigate the domain gap between different LiDAR scene flow datasets and show that multi-dataset training enables strong generalization. Due to a lack of benchmarks and baselines in diverse domains, our experiments are limited to autonomous driving scenes. Future work should explore non-AV settings.
5 Conclusion
In this paper, we introduce UniFlow, a frustratingly simple approach that re-trains off-the-shelf LiDAR scene flow models with diverse data from multiple datasets. Although prior work in LiDAR-based 3D object detection and segmentation don’t seem to benefit from multi-dataset training, we posit that our proposed method works well “low-level” tasks such as motion estimation may be less sensitive to sensor configuration. Notably, our model establishes a new state-of-the-art Waymo and nuScenes, improving over prior work by 5.1% and 35.2% respectively. Although UniFlowis primarily trained on urban city driving scenarios, we demonstrate that it generalizes surprisingly well to highway truck driving without any dataset-specific fine-tuning. Our extensive analysis shows that such improvements are model-agnostic, suggesting that future scene flow methods should adopt a multi-dataset training strategy.
Appendix A Implementation Details
Models. We train and evaluate SSF and Flow using their official implementations without modification. As shown in Table 9, Flow4D benefits from a larger backbone when trained on the unified dataset. We denote this model as Flow4D-XL, which preserves the published architecture and increases all channel widths in the 4D Voxel Network by a uniform factor of three, with the Voxel Feature Encoders and Point Head adjusted accordingly to match the widened input/output dimensions. The full configuration is shown in Table 12. All dataset-specific Flow4D models use the original backbone, while all Flow4D models trained on the unified dataset use Flow4D-XL unless otherwise noted.
| Stage | STDB-P (Filters) | Pool & Up (Kernel, Stride) | Output shape |
| Input | – | – | |
| 1 | (48, 96) | Pool (2,2,2,1) | |
| (96, 96) | |||
| 2 | (96, 192) | Pool (2,2,2,1) | |
| (192, 192) | |||
| 3 | (192, 192) | Pool (2,2,2,1) | |
| (192, 192) | |||
| 4 | (192, 192) | Pool (2,2,1,1) | |
| (192, 192) | |||
| 5 | (192, 192) | Up (2,2,1,1) | |
| (192, 192) | |||
| 6 | (192, 192) | Up (2,2,2,1) | |
| (192, 192) | |||
| 7 | (192, 192) | Up (2,2,2,1) | |
| (192, 192) | |||
| 8 | (192, 192) | Up (2,2,2,1) | |
| (192, 192) | |||
| 9 | (96, 48) | – |
Augmentations. We apply two dataset-agnostic augmentations. Following Flow, we apply height jitter with probability 0.8 by translating the point cloud along the sensor’s z-axis using a uniform offset in the range of [0.5, 2.0] m. To improve cross-dataset density generalization, we also apply LiDAR beam dropout with probability 0.35, removing every other LiDAR beam for all Waymo and Argoverse 2 sweeps so that the effective density approximates a 32-beam sensor.
Training Settings. For the Argoverse 2 benchmark, we follow the standard evaluation protocol and report test-set performance according to the official public leaderboard to ensure fair comparison. For all TruckScenes experiments, baseline methods and their UniFlowvariants are trained under the same setups described in their respective original papers. UniFlow models adopt identical architectures unless otherwise noted, and all models are trained until convergence. We apply height augmentation to UniFlow and all dataset-specific Flow models to match the original training setup, and apply sparsity augmentation only to UniFlow models.
Appendix B Ablation on Synthetic Sparse LiDARs
We evaluate the performance of UniFlow models on synthetically sparsified LiDAR point clouds to investigate the impact of LiDAR sparsity on cross-domain generalization. We do not train any new models; all results are from models trained on the original datasets. Sparse datasets are created by removing every other beam from each LiDAR sweep. As shown in Table 13, the colored deltas indicate the change in performance relative to each model’s performance on the standard datasets (higher EPE indicates worse performance). Across datasets, degradation correlates strongly with dataset sparsity: models evaluated on Sparse nuScenes perform the worst due to its low (16-beam) density, while models evaluated on Sparse Waymo are affected the least due to its substantially higher point density.
Notably, UniFlow models are consistently more robust than dataset-specific baselines, exhibiting smaller performance drops across all datasets. This suggests that multi-dataset training provides useful regularization even under extreme sparsity.
| Method | Three-way EPE (cm) | Dynamic Bucket-Normalized | |||||||
| Mean | FD | FS | BS | Dynamic M | CAR | OTHER | PED | VRU | |
| Sparse AV2 | |||||||||
| SSF | 4.51 ↑0.44 | 10.79 ↑1.10 | 1.96 ↑0.21 | 0.80 ↑0.01 | 0.289 ↑0.041 | 0.183 ↑0.016 | 0.329 ↑0.061 | 0.336 ↑0.054 | 0.309 ↑0.032 |
| SSF (UniFlow, Ours) | 3.73 ↑0.24 | 9.29 ↑0.62 | 1.38 ↑0.11 | 0.53 ↑0.01 | 0.240 ↑0.018 | 0.160 ↑0.009 | 0.168 ↑0.003 | 0.328 ↑0.040 | 0.304 ↑0.019 |
| Flow4D | 3.99 ↑0.37 | 9.58 ↑1.03 | 1.70 ↑0.08 | 0.69 ↑0.00 | 0.209 ↑0.017 | 0.158 ↑0.012 | 0.183 ↑0.019 | 0.238 ↑0.017 | 0.255 ↑0.020 |
| Flow4D (UniFlow, Ours) | 3.44 ↑0.20 | 8.27 ↑0.99 | 1.39 ↑0.13 | 0.65 ↑0.02 | 0.181 ↑0.005 | 0.138 ↑0.005 | 0.151 ↑0.002 | 0.198 ↓0.005 | 0.237 ↑0.017 |
| Flow | 3.30 ↑0.26 | 8.00 ↑0.67 | 1.22 ↑0.09 | 0.67 ↑0.03 | 0.218 ↑0.042 | 0.149 ↑0.013 | 0.169 ↑0.008 | 0.230 ↑0.036 | 0.218 ↑0.005 |
| Flow (UniFlow, Ours) | 3.19 ↑0.17 | 7.53 ↑0.40 | 1.25 ↑0.07 | 0.77 ↑0.03 | 0.182 ↑0.010 | 0.141 ↑0.008 | 0.157 ↑0.002 | 0.219 ↑0.025 | 0.213 ↑0.004 |
| Sparse Waymo | |||||||||
| SSF | 2.45 ↑0.24 | 6.12 ↑0.50 | 0.92 ↑0.18 | 0.30 ↑0.05 | 0.217 ↓0.040 | 0.100 ↑0.003 | - | 0.342 ↓0.127 | 0.210 ↑0.003 |
| SSF (UniFlow, Ours) | 2.02 ↑0.17 | 5.22 ↑0.36 | 0.65 ↑0.10 | 0.18 ↑0.05 | 0.199 ↓0.042 | 0.085 ↑0.004 | - | 0.318 ↓0.139 | 0.193 ↑0.010 |
| Flow4D | 2.00 ↑0.18 | 4.92 ↑0.34 | 0.76 ↑0.15 | 0.33 ↑0.05 | 0.196 ↓0.018 | 0.076 ↑0.001 | - | 0.363 ↓0.059 | 0.150 ↑0.004 |
| Flow4D (UniFlow, Ours) | 1.66 ↑0.05 | 3.92 ↑0.09 | 0.73 ↑0.05 | 0.31 ↓0.02 | 0.143 ↓0.046 | 0.063 ↓0.001 | - | 0.249 ↓0.151 | 0.119 ↑0.015 |
| Flow | 1.53 ↑0.09 | 3.42 ↑0.12 | 0.68 ↑0.08 | 0.50 ↑0.09 | 0.146 ↓0.050 | 0.072 ↑0.002 | - | 0.257 ↓0.154 | 0.107 ↑0.000 |
| Flow (UniFlow, Ours) | 1.43 ↑0.05 | 3.12 ↑0.05 | 0.66 ↑0.05 | 0.52 ↑0.06 | 0.138 ↓0.050 | 0.066 ↑0.001 | - | 0.246 ↓0.157 | 0.102 ↑0.006 |
| Sparse nuScenes | |||||||||
| SSF | 3.96 ↑0.96 | 9.24 ↑2.69 | 2.18 ↑0.14 | 0.45 ↑0.05 | 0.298 ↑0.078 | 0.173 ↑0.031 | 0.223 ↑0.027 | 0.556 ↑0.158 | 0.241 ↑0.096 |
| SSF (UniFlow, Ours) | 2.94 ↑1.05 | 7.11 ↑3.01 | 1.56 ↑0.13 | 0.14 ↑0.01 | 0.222 ↑0.089 | 0.094 ↑0.013 | 0.165 ↑0.034 | 0.454 ↑0.225 | 0.173 ↑0.083 |
| Flow4D | 4.31 ↑0.86 | 10.41 ↑2.44 | 1.92 ↑0.07 | 0.61 ↑0.06 | 0.309 ↑0.078 | 0.227 ↑0.067 | 0.283 ↑0.042 | 0.439 ↑0.094 | 0.287 ↑0.056 |
| Flow4D (UniFlow, Ours) | 3.72 ↑0.71 | 9.31 ↑2.03 | 1.52 ↑0.05 | 0.31 ↑0.03 | 0.279 ↑0.083 | 0.140 ↑0.003 | 0.266 ↑0.047 | 0.430 ↑0.158 | 0.279 ↑0.122 |
| Flow | 2.76 ↑0.71 | 6.06 ↑1.87 | 1.46 ↑0.10 | 0.75 ↑0.16 | 0.255 ↑0.081 | 0.179 ↑0.047 | 0.217 ↑0.033 | 0.409 ↑0.148 | 0.214 ↑0.097 |
| Flow (UniFlow, Ours) | 2.25 ↑0.62 | 4.81 ↑1.63 | 1.43 ↑0.14 | 0.51 ↑0.09 | 0.206 ↑0.067 | 0.126 ↑0.028 | 0.170 ↑0.033 | 0.365 ↑0.133 | 0.165 ↑0.074 |
Appendix C Ablation on Velocity Augmentation
Our velocity-bucket analysis in Figure 2 suggests that the velocity domain gap between training datasets is highly correlated with cross-domain generalization. To test whether increasing the proportion of fast motion improves flow estimation on fast moving objects across datasets, we create a “fast” version of each dataset by down-sampling each dataset from 10Hz to 5Hz and add these sequences to the unified training set. This effectively doubles the speed of all dynamic objects and shifts the velocity distribution to the right (Figure 5). As shown in Table 14, this improves performance in the highest-speed bucket , but consistently degrades accuracy for all other velocity buckets. We attribute this to the aggressiveness of the augmentation: doubling the speed of every dynamic object produces unrealistic trajectories that no longer resemble any real object motion. Although the velocity domain gap significantly contributes to cross-domain generalization, naively down-sampling datasets to increase the proportion of fast moving objects does not improve performance. Future work should consider different strategies for effective velocity augmentation.
| Method | Dynamic Bucket-Normalized | Speed Bucket-Normalized | |||||||
| Dynamic M | CAR | OTHER | PED | VRU | [0,0.5] | (0.5,1.0] | (1.0,2.0] | (2.0,] | |
| Argoverse 2 | |||||||||
| Flow4D | 0.192 | 0.146 | 0.164 | 0.221 | 0.236 | 0.020 | 0.119 | 0.114 | 0.103 |
| Flow4D (UniFlow, Ours) | 0.176 | 0.133 | 0.149 | 0.203 | 0.220 | 0.018 | 0.107 | 0.108 | 0.091 |
| Flow4D (Fast-UniFlow, Ours) | 0.179 | 0.134 | 0.154 | 0.204 | 0.224 | 0.019 | 0.114 | 0.111 | 0.085 |
| Waymo | |||||||||
| Flow4D | 0.214 | 0.074 | - | 0.423 | 0.146 | 0.007 | 0.064 | 0.051 | 0.130 |
| Flow4D (UniFlow, Ours) | 0.191 | 0.064 | - | 0.400 | 0.110 | 0.006 | 0.053 | 0.042 | 0.128 |
| Flow4D (Fast-UniFlow, Ours) | 0.208 | 0.066 | - | 0.409 | 0.149 | 0.007 | 0.058 | 0.043 | 0.093 |
| nuScenes | |||||||||
| Flow4D | 0.230 | 0.160 | 0.241 | 0.345 | 0.176 | 0.013 | 0.100 | 0.109 | 0.370 |
| Flow4D (UniFlow, Ours) | 0.196 | 0.137 | 0.219 | 0.272 | 0.157 | 0.009 | 0.082 | 0.089 | 0.374 |
| Flow4D (Fast-UniFlow, Ours) | 0.187 | 0.126 | 0.217 | 0.263 | 0.143 | 0.009 | 0.090 | 0.101 | 0.306 |
| TruckScenes | |||||||||
| Flow4D | 0.456 | 0.176 | 0.351 | 0.885 | 0.413 | 0.054 | 0.192 | 0.171 | 0.091 |
| Flow4D (UniFlow, Ours) ZS | 0.281 | 0.088 | 0.277 | 0.530 | 0.230 | 0.034 | 0.097 | 0.119 | 0.389 |
| Flow4D (Fast-UniFlow, Ours)ZS | 0.281 | 0.079 | 0.264 | 0.546 | 0.235 | 0.034 | 0.112 | 0.106 | 0.323 |
Appendix D Performance Breakdown on TruckScenes
We present a detailed breakdown between three-way EPE and dynamic bucket-normalized EPE performance on TruckScenes to investigate the inconsistent trends found in the main paper. As demonstrated by Khatri et. al. [khatri2024can], three-way EPE directly measures absolute endpoint error and therefore scales with the magnitude of the ground-truth flow. As a result, underestimating large displacement vectors from fast moving objects in highway scenes incurs disproportionately high penalties, inflating three-way EPE for TruckScenes. In contrast, dynamic bucket-normalized EPE offers a more balanced assessment of model predictions by evaluating error relative to object speed. As shown in Table 15, UniFlow achieves strong zero-shot generalization compared to in-domain models across common-speed buckets (0 to 2.0 m/frame), indicating that UniFlow effectively models the motion patterns dominating most scenes. We find that UniFlow models perform worse than in-domain models in the high-speed regime ( m/frame), suggesting that future work should explicitly address high-speed scene flow.
| Method | Three-way EPE (cm) | Dynamic Bucket-Norm. | Speed Bucket-Norm. | ||||||||||
| Mean | FD | FS | BS | Dyn. M | CAR | OTHER | PED | VRU | [0,0.5] | (0.5,1.0] | (1.0,2.0] | (2.0,] | |
| SSF | 6.20 | 15.17 | 1.01 | 2.44 | 0.453 | 0.215 | 0.308 | 0.875 | 0.414 | 0.055 | 0.190 | 0.173 | 0.086 |
| SSF (UniFlow, Ours) | 35.23 | 103.72 | 1.69 | 0.27 | 0.435 | 0.149 | 0.455 | 0.669 | 0.466 | 0.036 | 0.148 | 0.285 | 0.637 |
| Flow4D | 16.14 | 44.87 | 1.71 | 1.85 | 0.456 | 0.176 | 0.351 | 0.885 | 0.413 | 0.027 | 0.249 | 0.157 | 0.079 |
| Flow4D (UniFlow, Ours) | 23.59 | 68.41 | 1.78 | 0.57 | 0.281 | 0.088 | 0.277 | 0.530 | 0.230 | 0.034 | 0.097 | 0.119 | 0.389 |
| Flow | 7.28 | 16.26 | 1.36 | 4.52 | 0.402 | 0.196 | 0.400 | 0.690 | 0.323 | 0.079 | 0.222 | 0.203 | 0.102 |
| Flow (UniFlow, Ours) | 15.76 | 45.76 | 1.02 | 1.32 | 0.283 | 0.101 | 0.293 | 0.513 | 0.226 | 0.039 | 0.094 | 0.127 | 0.429 |
Appendix E Visualizing Failure Cases
Scene 1 |
Scene 2 |
(a) Ground truth
(b) UniFlow
(c) Camera View
Figure 6 illustrates two representative failure modes of UniFlow models. In the first example, adverse weather introduces dense rain streaks that create spurious LiDAR returns, leading the model to produce localized artifacts. In the second example, extreme sparsity at long range causes the model not estimate flow for an object located nearly 100m away.
Appendix F Supplementary Video
We include a supplementary video on our project page, where we visualize scene flow predictions on a challenging rainy sequence from TruckScenes. As shown in the video, Flow frequently produces artifacts on rain streaks and background points, which become especially pronounced during occlusions. Flow also predicts inconsistent flow vectors on dynamic objects. In contrast, Flow (UniFlow) yields significantly more stable and coherent motion fields.